OpenStudy (anonymous):

A refrigerator repairmen charges a fixed amount for a service call in addition to his hourly rate. If a two-hour repair costs $50 and a four-hour repair costs $74, what is the repairman's hourly rate?

OpenStudy (anonymous):

Let x= fixed rate Let y=Hourly rate x+2y=50 x+4y=74

OpenStudy (anonymous):

Can you do the rest?

OpenStudy (anonymous):

@jocelynna97

OpenStudy (anonymous):

yes wouldn't i just from then use the substitution method? @Dido525

OpenStudy (anonymous):

Remember, each shirt costs $15.

OpenStudy (anonymous):

Or elimination which I find easier. You can use substitution if you want.

OpenStudy (anonymous):

@geoffb lol this is a different probelm...

OpenStudy (anonymous):

So x=50-2y

OpenStudy (anonymous):

Substitute that for x in the second one.

OpenStudy (anonymous):

I am sure you can figure it out now.

OpenStudy (anonymous):

I know. But my point is that it's essentially the exact same problem, so I don't know why you need help again—especially since you didn't show your work or where you're getting stuck.

OpenStudy (anonymous):

@geoffb i'm just not sure on how to set up the problem when i set it up then i know how to do the rest & clearly i wouldn't have posted this question if i was understanding what i need to do i thought this website was for helping and not judging

OpenStudy (anonymous):

@Dido525 Thank you so so so much for the help I appreciate it :)

OpenStudy (anonymous):

Welcome! :) .

OpenStudy (anonymous):

Who's judging? I thought the answer @ChmE gave in the other post was more than helpful, and would adequately prepare you for this one (since it's the exact same question).

OpenStudy (anonymous):

@geoffb I'm not great with word problems, obviously this isnt going anywhere thanks anyways but i understand which in the end is why i came here...